OpenStudy (anonymous):

Since 1953, when Frank graduated from high school, he has gained two pounds every year. In 1983 he was 40 percent heavier than in 1953. What percent of his 1998 weight was his 1983 weight?

OpenStudy (anonymous):

tips?

OpenStudy (anonymous):

From 1953 to 1983: 60 pounds heavier 40% = 60lbs 60/40 x 100 = 150 pounds in 1953 In 1998: 240 pounds.

OpenStudy (anonymous):

150/240 x 100 = 62.5%

OpenStudy (anonymous):

k

OpenStudy (anonymous):

In 1983, he weighted 210 pounds

OpenStudy (anonymous):

In 1998, he weighed 240 pounds

OpenStudy (anonymous):

In 1983, he weighted 150, your answer is wrong.

OpenStudy (anonymous):

nooo...

OpenStudy (anonymous):

2 pounds per year from 1953 to 1983. That's a 30 years difference. 30 x 2 pounds per year = 60 pounds. The 60 pounds is 40% 60/40 x 100 = 150

OpenStudy (anonymous):

60/40 x 100 = 150 pounds in 1953

OpenStudy (anonymous):

In 1998: 240 pounds.

OpenStudy (anonymous):

What percent of his 1998 weight was his 1983 weight?

OpenStudy (anonymous):

OH!! I'm sorry the question is to find from 1983, not 1953.

OpenStudy (anonymous):

87.5

OpenStudy (anonymous):

percent

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

just saying